OpenStudy (anonymous):

NEED SOMEONE TO SOLVE TO CHECK MY WORK second derivative of x^2 - y^2=16 dy/dx URGENT PLEASE THANK YOU

OpenStudy (anonymous):

\[2x-2yy'=0\] solve for \(y'\)

OpenStudy (anonymous):

oooh second derivative first is \[y'=\frac{x}{y}\] so you need the derivative of that one use the quotient rule

OpenStudy (anonymous):

\[y''=\frac{y-xy'}{y^2}\] then replace \(y'\) by \(\frac{x}{y}\)

OpenStudy (loser66):

dan, Question: how satelite73 has that y ' ?

OpenStudy (anonymous):

thank you!!! exactly what I got

OpenStudy (anonymous):

is it ok to simplify to 1-x dy/dx over y?

OpenStudy (jhannybean):

\[\large f(x)=x^2-y^2=16\]\[\large f'(x): 2x -2yy'=0\]\[\large f'(x): y'=\frac{x}{y}\]\[\large f''(x):y''=\frac{(1)y-xy'}{y^2}\]\[\large f''(x): y''=\frac{y-\frac{x}{y}}{y^2}\]\[\large f''(x):y''= \frac{\frac{y^2-x}{y}}{y^2}\]

OpenStudy (dan815):

x^2 - y^2=16dy/dx (x^2 - y^2)/16= dy/dx

OpenStudy (jhannybean):

\[\large \frac{y^2-x}{y^3}= \frac{1}{y}-\frac{x}{y^3}\]

OpenStudy (loser66):

yes, i'm with you dan, how about dy/dx at the last place?

OpenStudy (dan815):

d^2y/dx^2=dy/dx[(x^2 - y^2)/16]

OpenStudy (anonymous):

@Loser66 $$x^2-y^2=C\\\frac{d}{dx}[x^2-y^2]=0\\2x-2y\frac{dy}{dx}=0\\y\frac{dy}{dx}=x\\\frac{dy}{dx}=\frac{x}y\\$$

OpenStudy (loser66):

look at the problem, it has dy/dx after 16

OpenStudy (dan815):

i think that shudnt be there

OpenStudy (jhannybean):

@Loser66 the 2's get canceled out. and you're just left with x's and y's.

OpenStudy (dan815):

i did it with that dy/dx too lol

OpenStudy (jhannybean):

dy/dx means y w.r.t x. Or known as y prime= y'

OpenStudy (loser66):

without it, it's quite easy to understand

OpenStudy (jhannybean):

It's just a different format.

OpenStudy (anonymous):

I'm not really sure what he's asking tbh

OpenStudy (dan815):

ya i thought he was asking (x^2 - y^2)/16=dy/dx

OpenStudy (jhannybean):

differentials?

OpenStudy (dan815):

but i guess he just meant to say find d^2/dx^2=?

OpenStudy (jhannybean):

I solved it the way you would solve a second derivative....

OpenStudy (dan815):

d^2y/dx^2=?

OpenStudy (jhannybean):

Dan, how about just y''. lol.

OpenStudy (loser66):

ok, there you go, shut up

OpenStudy (dan815):

we are going to get banned loser u and me!

OpenStudy (loser66):

The question is unclear. I wonder whether it is x^2-y^2 =16y' ?

OpenStudy (dan815):

loser go read a book

OpenStudy (loser66):

if it is just x^2 -y^2 =16. ha!! @dan815 yes sir

OpenStudy (dan815):

give it here?